OpenStudy (anonymous):

if y=81+17x(square) find dy/dx?

OpenStudy (lalaly):

dy/dx = 34x

OpenStudy (lalaly):

its 17x^2 so 34x

OpenStudy (anonymous):

34x=dy/dx

OpenStudy (anonymous):

dy/dx of any Constant is always 0 so leave 81

OpenStudy (anonymous):

hey thankx bt how did u solve this equation.

OpenStudy (anonymous):

okay look To differentiate there are some rules there is a power rule which is for variables with powers We multiply the variable with its power and decrease the power by one lilke X^2=2*X^2-1

OpenStudy (anonymous):

ok thankx now i got the equation

OpenStudy (anonymous):

Yw so a medal will work :p

OpenStudy (anonymous):

sure